OpenStudy (anonymous):

integrate this too. please. int. sqr (1 + cos x dx) sorry if i'm asking too many questions. i have too many assignments and i'm confused on most integrations.

OpenStudy (anonymous):

\[\int\limits \sqrt{1 + cosx} dx\]

OpenStudy (anonymous):

zed, can u explain a little bit more. just a little bit more. thank you.

OpenStudy (anonymous):

sorry i forgot we were doing integrals (I was doing derivatives). Use double angle formula: cos(2θ) = cos²(θ) - sin²(θ) = cos²(θ) - (1 - cos²(θ)) = 2∙cos²(θ) - 1 with θ = (x/2) cos(x) = 2∙cos²(x/2) - 1 Sub this in to make the integral easier.

OpenStudy (lollylau):

have you ever tried to use the fundamental theorem of calculus?

OpenStudy (anonymous):

but i still don't understand. i'm sorry, i know im being stupid here, but how come i'm supposed to use a double angle formula in this problem?

OpenStudy (anonymous):

i used this here because it will get rid of the sqrt sign in the integral making it easier to integrate.

OpenStudy (anonymous):

ok. i see. thank you again and again. i'll try it out. :D